OpenStudy (jennychan12):

Light of wavelength 550 nm falls on a slit that is 3.50 x 10^-3 mm wide. How far from the central maximum will the first diffraction maximum fringe be if the screen is 10.0 m away?

OpenStudy (jennychan12):

Light of wavelength 550 nm falls on a slit that is 3.50 x 10^-3 mm wide. How far from the central maximum will the first diffraction maximum fringe be if the screen is 10.0 m away? This is what i have so far: λ = 550 * 10^{-9}\] m. D = 3.50*10^-6 m. d = 10 m. m = 1 2x = ? Is that right?

OpenStudy (anonymous):

you'll need find x when n=1